OpenStudy (anonymous):

What are the exact solutions of x2 - 5x - 1 = 0? Answer x = the quantity of negative 5 plus or minus the square root of 21 all over 2 x = the quantity of 5 plus or minus the square root of 29 all over 2 x = the quantity of 5 plus or minus the square root of 21 all over 2 x = the quantity of negative 5 plus or minus the square root of 29 all over 2

OpenStudy (anonymous):

second answer. \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]Negative b rules out all but the middle two, and since c is negative the discriminant is \[5^{2}-4(-1)(1)=25+4=29\]